The locus of the point z satisfying the condition arg (Z-1)/(Z+1)=Pi/3

Grade:12

1 Answers

Ravi
9 years ago
Transform the arg [(z-1)/(z+1)]=pi/3 into (z-1)/(z+1)=tan pi/3

Solve this eqn for z.. It will give a linear eqn. solution.
.........................................................................................................................................